A Bibliography of Publications in American

Mathematical Monthly: 1950–1959

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.04

Title word cross-reference

16 [8]. 61 [8]. Amer [8]. approximation [2]. associated [3]. best [2]. connected [6]. cores [7]. curve [3]. double [1]. Fan [8]. functions [2, 7].

G [8]. geometrically [1]. implicants [7]. inequality [4, 8]. integers [6]. integral [4, 8]. Lorentz [8]. Math [8]. Monthly [8]. MR [8]. nonassociative [5]. Note [8]. number [5]. polynomial [2]. prime [7]. series [1]. summed [1]. tetrahedron [3]. Theorems [5]. theory [5]. topology [6]. truth [7].

References

Tornheim:1950:DSS

[1] Leonhard Tornheim. A double series summed geometrically. American Mathematical Monthly, 57(??):535–538, ???? 1950. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Olds:1950:BPA

[2] C. D. Olds. The best polynomial approximation of functions. American Mathematical Monthly, 57(??):617–621, ???? 1950. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Thebault:1950:CAT

[3] Victor Thebault. A curve associated with a tetrahedron. American Mathematical Monthly, 57(??):663–667, ???? 1950. CO-DEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Fan:1954:II

[4] K. Fan and G. G. Lorentz. An integral inequality. American Mathematical Monthly, 61(??):626–631, ???? 1954. See note [8].

Minc:1959:TNN

[5] H. Minc. Theorems on nonassociative number theory. American Mathematical Monthly, 66(??):486–488, ???? 1959. CO-DEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Golomb:1959:CTI

[6] S. W. Golomb. A connected topology for the integers. American Mathematical Monthly, 66(??):663–665, ???? 1959. CO-DEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Quine:1959:CPI

[7] W. V. Quine. On cores and prime implicants of truth functions. American Mathematical Monthly, 66(??):755–760, ???? 1959. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pecaric:1981:NII

[8] Josip E. Pečarić. Note on: “An integral inequality” [Amer. Math. Monthly 61 (1954), 626–631, MR 16, 342] by K. Fan and G. G. Lorentz. American Mathematical Monthly, 88(4):275–276, ???? 1981. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [4].